the perimeter oa a cross section or end of a "two-by-four" piece of lumber is 10 in. The lengt is 2 in. more than the width. Find the actual dimensions of the cross section of a two-by-four.
Answer by stanbon(75887)

The perimeter of a cross section or end of a "two-by-four" piece of lumber is 10 in. The length is 2 in. more than the width.
Find the actual dimensions of the cross section of a two-by-four.
-------------------------
Perimeter = 2(W + L)
10 = 2(W + (W+2))
5 = 2W+2
2W = 3
Width = 1.5"
Length = 1.5+2 = 3.5"
==========================
Cheers,
Stan H.